Perform the calculation (8i)3-(8i)2+8i-1. Give your answer in Cartesian form z=x+iy.
Answer #1

given (8153–18172+81-1. = 83;3- 8?ci)?+81-1 = 512l-i)-642-1)+81-1 =-5121 +64 +81-1 - 64-1-5121+81 = ô 3 -soui : cartesian for